# The Impact of Unstated Norms in Bias Analysis of Language Models

### v4

@@ 3.3 Fine-Tuning Experimental Setup
The LLMs considered in this set of experiments are drawn from the RoBERTa Liu et al. (2019), OPT Zhang et al. (2022), and Llama-2/3 Touvron et al. (2023), and Mistral Jiang et al. (2023) families of models. Specifically, we consider RoBERTa 125M and 355M, OPT 125M, 350M, 1.3B, and 6.7B, Llama-2 7B and 13B, Llama-3 8B, and Mistral 7B. Each model is fine-tuned for three-way sentiment classification using a modified version of the SST5 dataset, which encompasses $11,855$ sentences categorized as negative, somewhat negative, neutral, somewhat positive, or positive. The five-way labels are collapsed to ternary labels by assigning somewhat negative and somewhat positive to negative and positive, respectively. OPT 125M and 350M and RoBERTa 125M and 355M are fully fine-tuned. Due to their size, the remaining models are fine-tuned with LoRA Hu et al. (2022) using a bottleneck dimension of $8$ on every non-embedding layer.
@@
During training, early stopping is applied based on validation loss. If no improvement in the loss is observed over a fixed number of steps, then training is stopped. An AdamW optimizer is used with default parameters, except for learning rate (LR) and weight decay Loshchilov and Hutter (2019). A hyper-parameter study was performed to select the best early stopping threshold and LR for all models. For fully fine-tuned models, weight decay was also optimized. The early stopping threshold was varied between five and seven steps. The LR was selected from $\{1\mathrm{e}{-3},3\mathrm{e}{-4},1\mathrm{e}{-4},3\mathrm{e}{-5},1\mathrm{e}{-5}\}$. Finally, weight decay, when tuned, was selected from $\{1\mathrm{e}{-3},1\mathrm{e}{-4},1\mathrm{e}{-5},1\mathrm{e}{-6}\}$. Table 1 displays the optimal hyper-parameters used to train the final models. For RoBERTa 125M and 355M and OPT 125M and 350M, $15$ training runs are performed and the five models with the highest accuracy on the SST5 test set are retained. For the larger models, due to resource constraints, five models in total are trained for each model type.
@@
RoBERTa-125M $7$ $1\mathrm{e}{-5}$ $1\mathrm{e}{-5}$
RoBERTa-355M $7$ $1\mathrm{e}{-5}$ $1\mathrm{e}{-5}$
OPT-125M $7$ $1\mathrm{e}{-5}$ $1\mathrm{e}{-5}$
OPT-350M $7$ $1\mathrm{e}{-5}$ $1\mathrm{e}{-3}$
OPT-1.3B $5$ $1\mathrm{e}{-4}$ $1\mathrm{e}{-4}$
OPT-6.7B $5$ $1\mathrm{e}{-4}$ $1\mathrm{e}{-4}$
OPT-13B $5$ $1\mathrm{e}{-4}$ $1\mathrm{e}{-4}$
Llama-2-7B $5$ $1\mathrm{e}{-4}$ $1\mathrm{e}{-4}$
Llama-2-13B $5$ $1\mathrm{e}{-4}$ $1\mathrm{e}{-4}$
Llama-3-8B $5$ $1\mathrm{e}{-4}$ $1\mathrm{e}{-3}$
Mistral-7B $5$ $3\mathrm{e}{-5}$ $1\mathrm{e}{-3}$
Table 1: Hyper-parameters for model fine-tuning.
To measure model performance disparities across races, each of the trained models performs inference on examples generated from the three datasets discussed in Sections 3.1.1-3.1.3 to predict their sentiment. Using these predictions, FPR gaps are computed for examples associated with the different racial groups. Training a set of models facilitates the computation of 95% CIs for the gaps, which are reported alongside the mean gaps.
3.4 Prompting Experimental Setup
Three prompting strategies are applied to predict sentiment. These are zero-shot prompts, 9-shot prompts with shots drawn from two sentiment analysis datasets, and zero-shot CoT prompts Kojima et al. (2024). For all prompting experiments, Hugging Face’s text-generation pipeline is used for the base models of OPT-6.7B, Llama-2-7B, Llama-3-8B, and Mistral-7B. While local model artifacts are used in the experiments, these models correspond to the Hugging Face identifiers facebook/opt-6.7b, meta-llama/Llama-2-7b-hf, meta-llama/Meta-Llama-3-8B, and mistralai/Mistral-7B-v0.1. Sampling is turned on, and a temperature of 0.8 is used for all generations, including reasoning traces. Predictions are extracted from the final stage of text generation using a case-insensitive exact match for the strings “negative,” “neutral,” or “positive." The first instances of such a match are taken as the predicted label. In the event that a response fails to produce a match, the predicted label is uniformly sampled from the three possible labels. In all but the reasoning generation stage of zero-shot CoT, models produce a maximum of three tokens in their response.
@@ 3.4 Prompting Experimental Setup
[…]
@@
The final prompting approach, zero-shot CoT, uses two prompt templates in sequence. Note that CoT prompting is not used for OPT, as the model has been shown to have limited reasoning capacity Liang et al. (2023). In the first step, the model is provided the text to classify and asked about the corresponding sentiment. The traditional “trigger” sentence “Let’s think step by step.” is used to encourage the model to generate reasoning prior to answering the question. Reasoning traces are capped at 64 tokens. The template appears below.
@@
For CoT prompting, inference batches are of size $4$ due to the larger computation required, whereas batch sizes of $16$ are applied in the previous settings. To quantify generation stochasticity, predictions for each example are generated five times.
4 Results
4.1 Fine-Tuning Results
The Negative- and Positive-Sentiment FPR gaps for the Amazon dataset are shown in Figure 1. For most models, the negative-sentiment FPR gap for White-associated text is significantly above zero at 95% confidence. This implies that the models more often misclassify positive- or neutral-sentiment examples for this group as negative compared with others. For large OPT, Llama-2 and Mistral LLMs, a similar but smaller elevation in this gap is observed for examples associated with African Americans and Asians. For the Positive-Sentiment FPR gap, a significant negative value is observed for all models. Recent models, Llama-3 and Mistral, exhibit some of the largest negative gaps. Combined with an elevated Negative-Sentiment FPR gap, this implies that the models tend to view examples from the White race in a negative light more often than other groups.
Figure 2 displays the measured gaps for the NS-Prompts dataset. Recall that all labels for this dataset are neutral. Thus, any non-neutral predictions are, by definition, incorrect. When considering RoBERTa and Llama-2 models, the identified gaps share similarities with the African-American group. That is, elevated Negative-Sentiment FPR gaps and Positive-Sentiment FPR gaps below zero. While the negative-sentiment FPR gaps for other models are near zero for White examples, all models produce negative and statistically significant Positive-Sentiment FPR gaps. This implies that neutral examples associated with White race are construed as positive at lower rates relative to other groups.
@@ 4.1 Fine-Tuning Results
Table 2: Model accuracy and standard deviation on the Amazon dataset for prompting experiments across model types. Bold numbers indicate the best accuracy achieved for each model.
The measurements in these results are surprising. However, the gaps observed for the White group are not believed to be reflections of true bias. Rather, they are likely an artifact of a mismatch between the template-based probing datasets that explicitly reference race to link membership and markedness in LLM pretraining data.
4.2 Prompt-Based Results
The results in Section 4.1 exhibit clear anomalies when measuring performance gaps using template-based probes. A natural question is whether such irregularities arise due to the task-specific fine-tuning step or represent an intrinsic quality of the LLMs. To further isolate the issue to LLM pretraining, prompting is used to perform sentiment classification for the Amazon dataset, shedding the need for fine-tuning. The experiments are limited to decoder-only models of sufficient size to ensure that classification performance adequately exceeds that of a random classifier. The accuracy of these approaches is generally lower than that achieved in the fine-tuning experiments, but remains reasonable.
Model Size Avg. Acc. Std. Dev.
RoBERTa 125M $\mathbf{0.635}$ $0.036$
350M $0.624$ $0.027$
OPT 125M $0.687$ $0.080$
350M $0.692$ $0.039$
1.3B $\mathbf{0.739}$ $0.020$
6.7B $0.737$ $0.014$
Llama-2 7B $0.513$ $0.089$
13B $\mathbf{0.647}$ $0.006$
Llama-3 8B $\mathbf{0.822}$ $0.035$
Mistral 7B $\mathbf{0.740}$ $0.005$
Tables 2 and 3 display the average classification accuracy of the prompting and fine-tuning approaches on the Amazon dataset. Generally, the accuracy of prompt-based classification is lower than the fine-tuning counterpart. This is especially true for the oldest model, OPT. The best performing method is the 9-shot prompt drawn from SST5 with an accuracy of $71.6$% using Llama-3-8B. As seen in Table 3, many fine-tuned models approach or outperform this accuracy. Nonetheless, as classifiers, the prompted LLMs perform well above a random model. Perhaps due to size of the models, reasoning in the form of zero-shot CoT does not significantly improve performance Wei et al. (2024).
As in Section 4.1, Negative- and Positive-Sentiment FPR gaps are computed for each LLM’s predictions. These gaps are exhibited in Figure 4. Due to the lower accuracy and generation volatility, the gap CIs are visibly wider than those in the fine-tuning experiments. Nonetheless, a clear and familiar pattern is seen in these results. Positive mean gaps in Negative-Sentiment FPR are present across nearly all examples for African American and White races. Similarly, negative mean gaps for Positive-Sentiment FPR are measured for both races in most settings. The consistency between these results and those of the fine-tuning experiments strongly suggests that the irregularities present in the template-based measurements are not the result of fine-tuning, but rather an expression of an intrinsic aspect of the LLMs.
Figure 4: Negative- and Positive-Sentiment FPR gaps as measured by the Amazon dataset with prompt-based classification. In the legend, ZS stands for zero-shot while SST5 and SemEval indicate 9-shot, few-shot prompts with examples drawn from those datsets.
5 Discussion
We observed an overall tendency of the models to classify White-associated text as exhibiting negative sentiment at a higher rate than other groups. The trends observed in the above results are consistent between model types, model size, template-based probing dataset, and even classification strategies. The overall agreement of the prompting and fine-tuning results indicates that the observed gaps are not linked to idiosyncrasies in the fine-tuning process but are, rather, more fundamental to the LLMs themselves. In addition, the models chosen for experimentation are base versions. That is, their predictions are not influenced by interceding alignment techniques Bai et al. (2022); Rafailov et al. (2023), which might otherwise obscure behavior learned during pretraining. Rather than implying an extant bias, we hypothesize that this phenomenon is due to an interaction between the structure of the templates used in the measurement of bias and LLM pretraining data, leading us to re-think the clarity of counterfactual bias analysis in this context.

### v5

@@ 3.3 Fine-Tuning Experimental Setup
The LLMs considered in this set of experiments are drawn from the RoBERTa (Liu et al., 2020), OPT (Zhang et al., 2022), Llama-2/3 (Touvron et al., 2023), and Mistral (Jiang et al., 2023) families of models. Specifically, RoBERTa 125M and 355M, OPT 125M, 350M, 1.3B, and 6.7B, Llama-2 7B and 13B, Llama-3 8B, and Mistral 7B are considered. Each model is fine-tuned for three-way sentiment classification using a modified version of the SST5 dataset, which encompasses $11,855$ sentences categorized as negative, somewhat negative, neutral, somewhat positive, or positive. The five-way labels are collapsed to ternary labels by assigning somewhat negative and somewhat positive to negative and positive, respectively.
OPT 125M and 350M and RoBERTa 125M and 355M are fully fine-tuned. Due to their size, the remaining models are fine-tuned with LoRA (Hu et al., 2022). Each model is trained five separate times with different random seeds. Detailed hyperparameter settings for fine-tuning are included in Appendix A. To measure model performance disparities across races, each of the trained models performs inference on examples generated from the three datasets discussed in Sections 3.1.1-3.1.3 to predict their sentiment. Using these predictions, FPR gaps are computed for examples associated with the different racial groups. Training a set of models facilitates the computation of 95% CIs for the gaps, which are reported alongside the mean gaps.
3.4 Prompting Experimental Setup
Three prompting strategies are applied to predict sentiment. These are zero-shot prompts, 9-shot prompts with shots drawn from two sentiment analysis datasets, and zero-shot CoT prompts (Kojima et al., 2024). For all prompting experiments, Hugging Face’s text-generation pipeline is used with OPT-6.7B, Llama-2-7B, Llama-3-8B, Mistral-7B, Gemma-7B Instruct (Gemma et al., 2024), and Qwen-2.5-7B Instruct (Qwen et al., 2025). These models correspond to the Hugging Face identifiers facebook/opt-6.7b, meta-llama/Llama-2-7b-hf, meta-llama/Meta-Llama-3-8B, mistralai/Mistral-7B-v0.1, google/gemma-7b-it, and Qwen/Qwen2.5-7B-Instruct. Sampling is turned on, and a temperature of $0.8$ is used for all generations, including reasoning traces. Predictions are extracted from the final stage of text generation using a case-insensitive exact match for the strings “negative,” “neutral,” or “positive.” The first match is taken as the predicted label. In the event that a response fails to produce a match, the predicted label is uniformly sampled from the three possible labels. In all but the reasoning generation stage of zero-shot CoT, models produce a maximum of three tokens in their response.
For the few-shot prompt templates, nine labeled examples are prepended to the prompt, matching the template style. Two distinct experiments are conducted with labeled demonstrations drawn from either the SST5 or SemEval (Mohammad et al., 2018) datasets. For SST5, labels are collapsed as described in Section 3.3. The SemEval polarities are condensed via the mapping {Negative: [-3, -2], Neutral: [-1, 0, 1], Positive: [2, 3]}. In both cases, to avoid any few-shot bias (Gupta et al., 2024), demonstrations are balanced between negative, neutral, and positive (3 each), but order is random. Demonstrations are constant across models, but are resampled across the five prediction runs of each experiment. For all prompts, random seeds for shot selection and text generation are set to $\{2024$, $2025$, $2026$, $2027$, and 2028}2028\} across the five runs.
Zero-shot CoT uses two sequential prompt templates. CoT prompting is not applied to OPT, as the model has limited reasoning capacity (Liang et al., 2023). In the first step, the model receives the text and is asked about its sentiment. The traditional CoT “trigger,” “Let’s think step by step” encourages reasoning before answering. Reasoning traces are capped at $64$ tokens. To quantify generation stochasticity, each example is predicted five times. All prompt templates across strategies and other settings appear in Appendix B.
4 Results
4.1 Fine-Tuning Results
The Negative- and Positive-Sentiment FPR gaps for the Amazon dataset are shown in Figure 1. For most models, the Negative-Sentiment FPR gap for White-associated text is significantly above zero at 95% confidence. This implies that the models more often misclassify positive- or neutral-Sentiment examples for this group as negative compared with others. For large OPT, Llama-2 and Mistral LLMs, a similar but smaller elevation in this gap is observed for examples associated with African Americans and Asians. For the Positive-Sentiment FPR gap, a significant negative value is observed for all models. More recent models, Llama-3 and Mistral, exhibit some of the largest negative gaps. Combined with an elevated Negative-Sentiment FPR gap, this implies that the models tend to view examples from the White group in a negative light more often than other groups.
Figure 2 displays the measured gaps for the NS-Prompts dataset. Recall that all labels for this dataset are neutral. Thus, any non-neutral predictions are, by construction, incorrect. When considering RoBERTa and Llama-2 models, the identified gaps share similarities with the African-American group. That is, elevated Negative-Sentiment FPR gaps and Positive-Sentiment FPR gaps below zero. While the Negative-Sentiment FPR gaps for other models are near zero for White examples, all models produce negative and statistically significant Positive-Sentiment FPR gaps. This implies that neutral examples associated with White race are construed as positive at much lower rates relative to other groups.
The measurements in these results are surprising. However, as discussed in detail in Section 5 below, the gaps observed for the White group are not believed to be reflections of true bias. Rather, we conjecture that they are an artifact of a mismatch between the template-based probing datasets that explicitly reference race in the text of samples to link membership and the presence of markedness in LLM pretraining data.
4.2 Prompt-Based Results
The results in Section 4.1 exhibit clear anomalies when measuring performance gaps using template-based probes. A natural question is whether such irregularities arise due to the task-specific fine-tuning step or represent an intrinsic quality of the LLMs. To further isolate the issue to LLM pretraining, prompting is used to perform sentiment classification for the Amazon dataset, shedding the need for fine-tuning. The experiments are limited to decoder-only models of sufficient size to ensure that classification performance adequately exceeds that of a random classifier.
The average classification accuracy of the prompting and fine-tuning approaches on the Amazon dataset is reported in Appendix C. Generally, the accuracy of prompt-based classification is lower than the fine-tuning counterpart. This is especially true for the oldest model, OPT. However, newer models still provide good performance through prompting. Notably, Qwen-2.5 produces very strong classification accuracy with a 9-shot prompt drawn from SemEval at $92.3\%$. Other models also perform fairly well using few-shot prompting. Regardless, as classifiers, all prompted LLMs perform well above a random model. Perhaps due to model size or limited reasoning tokens, zero-shot CoT does not significantly improve performance (Wei et al., 2024).
As in Section 4.1, Negative- and Positive-Sentiment FPR gaps are computed for each LLM’s predictions. These gaps are exhibited in Figure 4. Due to the lower accuracy and generation volatility, the gap CIs are visibly wider than those of the fine-tuning experiments. Nonetheless, a clear and familiar pattern is seen in these results. Positive mean gaps in Negative-Sentiment FPR are present across a majority of examples for African American and White races. Similarly, negative mean gaps for Positive-Sentiment FPR are measured for both races in most settings. The consistency between these results and those of the fine-tuning experiments strongly suggests that the irregularities present in the template-based measurements are not the result of the fine-tuning stage, but are, rather, an expression of an intrinsic aspect of the LLMs themselves.
Figure 4: Negative- and Positive-Sentiment FPR gaps as measured by the Amazon dataset with prompt-based classification. In the legend, model names without a suffix indicate zero-shot prompting. SST5 and SE indicate 9-shot prompts with examples drawn from the SST5 and SemEval datasets, respectively.
5 Discussion
Across the experiments an overall tendency of the models to classify White-associated text as exhibiting more negative sentiment at a higher rate than other groups is observed. The trends in the results above are consistent between model type, model size, template-based probing dataset, and even classification strategy. The overall agreement of the prompting and fine-tuning results indicates that the observed gaps are not linked to idiosyncrasies in the fine-tuning process but are, rather, more fundamental to the LLMs themselves and the design of the template-based probes. In addition, the models chosen for experimentation are base or instruction-tuned versions. That is, their predictions are not influenced by interceding alignment techniques (Bai et al., 2022; Rafailov et al., 2023), which might otherwise obscure behavior learned during pretraining. Rather than implying an extant bias, we hypothesize below that this phenomenon is due to an interaction between the structure of the templates used in the measurement of bias and markedness in LLM pretraining data. Regardless of the underlying cause, these observations should lead us to re-think the clarity of counterfactual bias analysis in this context.
5.1 Markedness and Template-Based Probes
English pretraining data for LLMs is dominated by text drawn from areas where the racial majority is White (Bender et al., 2021; Navigli et al., 2023). Several studies have confirmed that markedness is widespread in internet data, with White race and male gender constituting the unmarked defaults (Wolfe and Caliskan, 2022b; Bailey et al., 2022). Furthermore, it has been shown that models, and LLMs in particular, trained on web data reflect these markedness characteristics (Bender et al., 2021; Wolfe and Caliskan, 2022a; b; Cheng et al., 2023). On the other hand, in templates commonly used for bias quantification, race is explicitly mentioned to establish group membership. As such, template-based text that explicitly establishes that the subject is “White” essentially constitute out-of-domain examples (Blodgett et al., 2021; Dressler, 1985). Such a mismatch likely influences model predictions.
We hypothesize that the disparities observed in Section 4 associated with the White group are due to the prevalence of markedness in LLM pretraining text. A key assumption underlying unmarked representations is that humans are adept at recognizing unstated implications in text. LLMs trained solely on unstructured next-token prediction, which underpins almost all modern LLM pretraining, may lack the ability to perceive such implications, resulting in surprising behavior. Using templates that represent group membership through explicit description likely makes certain text appear uncommon for traditionally unmarked groups. As such, these templates may lead to artificially elevated error rates in LLMs, skewing bias measurements in unpredictable ways and clouding the lens provided by datasets of this structure. To this end, Appendix F provides an extended set of results showing that similar irregularities arise when considering unmarked groups for sexuality and gender, providing additional supporting evidence for this hypothesis.
Including datasets that explicitly correct for markedness in LLM pretraining could better align template-based text. Appendix D suggests that more recent LLMs, trained on larger multilingual datasets, show improvements in measured gap sizes. Both Llama-3-8B and Mistral-7B have the smallest difference between the most positive and negative gaps for Negative-Sentiment FPR, averaged over the three datasets. Llama-3-8B also produces the lowest average difference for Positive-Sentiment FPR. Given that White-group gaps often rank among the extremes, this suggests newer models may be less affected by markedness.
The studies and results above, and in the appendix, suggest that markedness may indeed play a role in the experimental observations of this work. However, empirical validation of the conjecture that markedness is the mechanism introducing the effects observed in Section 4 likely requires that an LLM be trained from scratch using a dataset with “consistent” marking. Thereafter, the gaps across demographics would be re-assessed using the same template-based probes. The size and availability of LLM pretraining data make constructing such a dataset quite challenging. Moreover, the computational resources required to properly pretrain an LLM are substantial. As such, this valuable study is deferred to future work.
6 Conclusions and Future Work
This paper presents unexpected, and likely flawed, bias measurements related to race when using template-based bias probes. The measurements remain consistent across a number of different experimental settings and varied datasets. Rather than indicating genuine social bias in the LLMs, we conjecture that these outliers stem from a misalignment between template-based bias probes and LLM pretraining data due to markedness. Regardless of the underlying cause, these findings highlight the need to consider the impact that the use of bias probes relying on marked text has on the measurement of bias. In this case, such probes produce largely misleading results.
Assuming that linguistic markedness contributes significantly to the measurements in this work, which requires further investigation to confirm, several avenues for mitigating such effects when using template-based probes are worth exploring. LLMs trained on a more global representation of text, where majority demographics differ and marked groups vary, could improve the robustness of such models when encountering explicit demographic descriptors. Another approach is to strive to control for the impact of markedness by designing evaluation setups that test both unmarked and explicitly marked versions of the same text (for example, comparing “a CEO” and “a White CEO”) or by using neutral placeholders like [RACE] to isolate the impact of demographic terms and potentially correct for the flaws identified in this work. There are, however, challenges with this approach. One needs to identify which groups are considered unmarked from the “perspective” of the LLM, requiring detailed knowledge of the underlying pretraining data. In addition, the unassociated text is unlikely to solely represent the unmarked group, but rather a mix of representations. Ideally, artificial injection of demographic information would not be required. For example, the studies of Seyyed-Kalantari et al. (2020) and Sap et al. (2019) establish group membership through meta-data, self-identification, or classification techniques rather than explicitly in text. These methods avoid the out-of-domain nature of template-based examples of the kind studied here and do not see the unnatural patterns we observed. Future work will design experiments to validate the misalignment due to markedness conjecture and construct straightforward ways to mitigate such issues in LLMs.
Appendix A Fine-Tuning Hyperparameters
@@
For completeness, we provide the full details of the hyperparameter tuning process used in the fine-tuning experiments. During fine-tuning, early stopping is applied based on validation loss. If no improvement in the loss is observed over a fixed number of steps, then training is stopped. An AdamW optimizer is used with default parameters, except for learning rate (LR) and weight decay (Loshchilov and Hutter, 2019). A hyperparameter study was performed to select the best early stopping threshold and LR for all models. For fully fine-tuned models, weight decay was also optimized.
The early stopping threshold was varied between five and seven steps. The learning rate (LR) was chosen from {1e-3, 3e-4, 1e-4, 3e-5, 1e-5}, and weight decay, when tuned, was selected from {1e-3, 1e-4, 1e-5, 1e-6}. For larger models, LoRA fine-tuning was applied with the rank parameter 8 on every non-embedding layer.
For RoBERTa 125M and 355M and OPT 125M and 350M, 15 training runs were performed, and the five models with the highest accuracy on the SST5 test set were retained. For the larger models, due to resource constraints, five models in total were trained for each model type. Table 1 summarizes the optimal hyperparameters selected for each model during fine-tuning.
Table 1: Hyperparameters used for model fine-tuning.
@@
RoBERTa-125M 77 $1\mathrm{e}{-5}$ $1\mathrm{e}{-5}$
RoBERTa-355M 77 $1\mathrm{e}{-5}$ $1\mathrm{e}{-5}$
OPT-125M 77 $1\mathrm{e}{-5}$ $1\mathrm{e}{-5}$
OPT-350M 77 $1\mathrm{e}{-5}$ $1\mathrm{e}{-3}$
OPT-1.3B 55 $1\mathrm{e}{-4}$ $1\mathrm{e}{-4}$
OPT-6.7B 55 $1\mathrm{e}{-4}$ $1\mathrm{e}{-4}$
OPT-13B 55 $1\mathrm{e}{-4}$ $1\mathrm{e}{-4}$
Llama-2-7B 55 $1\mathrm{e}{-4}$ $1\mathrm{e}{-4}$
Llama-2-13B 55 $1\mathrm{e}{-4}$ $1\mathrm{e}{-4}$
Llama-3-8B 55 $1\mathrm{e}{-4}$ $1\mathrm{e}{-3}$
Mistral-7B 55 $3\mathrm{e}{-5}$ $1\mathrm{e}{-3}$
Appendix B Prompt Templates and Other Details
This section includes the templates used in the prompting approach. Each subsection corresponds to a different template. For CoT prompting, inference batches are limited to size $4$ due to higher computational demands, whereas batch sizes of $16$ are used in other settings.
B.1 Zero-Shot Prompt Template
@@
B.2 Few-Shot Prompt
Below is the few-shot template. For the few-shot prompt templates, nine labeled examples are prepended to the prompt, following the template style. The models begin generation at [LM Generation].
@@ B.2 Few-Shot Prompt
B.3 Zero-Shot CoT Prompt
Zero-shot CoT uses two prompt templates in sequence. In the first step, the model is provided the text to classify and asked about the corresponding sentiment. The traditional “trigger” sentence “Let’s think step by step” is used to encourage the model to generate reasoning prior to answering the question. The template appears below.
@@ B.3 Zero-Shot CoT Prompt
Model Size Mean Accuracy Standard Deviation
RoBERTa 125M 0.635 0.036
350M 0.624 0.027
OPT 125M 0.687 0.080
350M 0.692 0.039
1.3B 0.739 0.020
6.7B 0.737 0.014
Llama-2 7B 0.513 0.089
13B 0.647 0.006
Llama-3 8B 0.822 0.035
Mistral 7B 0.740 0.005
Appendix C Fine-tuning and Prompting Accuracy
Tables 2 and 3 present the average classification accuracy and standard deviation for the fine-tuning and prompting approaches on the Amazon dataset, respectively. Generally, prompt-based classification accuracy is lower than that of fine-tuning. There is also notable variability in model classification accuracy on the template-based probe dataset as a whole, with newer models tending to produce better performance. However, the trends associated with the measured FPR gaps, especially for the White group, are largely consistent, regardless of these variations. Do to this consistency, the differences in model accuracy are unlikely to be a primary contributor to the observed behavior.
Table 3: Model accuracy and standard deviation (in parentheses) on the Amazon dataset for prompting experiments across model types. Bold numbers indicate the best accuracy achieved for each model.
@@ Appendix C Fine-tuning and Prompting Accuracy
Gemma-7B $0.830$ ($0.001$) $0.777$ ($0.003$) $\mathbf{0.854}$ ($0.020$) $0.804$ ($0.031$)
Qwen-2.5-7B $0.899$ ($0.001$) $0.823$ ($0.002$) $\mathbf{0.923}$ ($0.016$) $0.906$ ($0.001$)